susan Mariano invested part of her $27000 bonus in a fund that paid an 11% profit and invested the rest in stock that suffered a 3% loss. Find the amount of each investment if her overall profit was $1990
Answer by checkley77(12844)

.11X-.03(27,000-X)=1,990
.11X-810+.03X=1,990
.14X=1,990+810
.14X=2,800
X=2,800/.14
X=$20,000 amount invested @ 11%
27,000-20,000=$7,000 invested @ 3%
Proof:
.11*20,000-.03*7,000=1,990
2,200-210=1,990
1,990=1,990